OpenStudy (anonymous):

The common difference in an arithmetic sequence is –2 and the first term is 47. What is the 29th term? A. –9 B. –11 C. 11 D. 18

sam (.sam.):

Figure out the formula to find nth term

OpenStudy (anonymous):

how do I do that?

OpenStudy (texaschic101):

here is the formula : an = a1 + (n-1)d a1 = your first term n = number of the term to find d = common difference lets begin... a29 = 47 + (29 - 1)(-2) a29 = 47 + (28)(-2) a29 = 47 + (-56) a29 = 47 - 56 a29 = -9 any questions ?

OpenStudy (texaschic101):

this might help... http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

okay ; will this be the formula for all the sequences?

OpenStudy (texaschic101):

this is the formula for an arithmetic sequence and finding a certain term....not the sum and not geometric sequences

OpenStudy (anonymous):

okay thank you

OpenStudy (texaschic101):

sure thing :)